OpenStudy (anonymous):

Solve the following equation: 7y - 3 ≤ 10y + 9

OpenStudy (anonymous):



OpenStudy (a1234):

7y - 3 + 3 <= 10y + 9 + 3 7y <= 10y + 12 -3y <= 12 y <= -4

OpenStudy (a1234):

Sorry...it's y >= 4, forgot to change the sign.

OpenStudy (a1234):

*-4